Find the greatest and smallest values that the function f(x,y) = 4xy takes on the ellipse x^2 + 2^y2 = 1 by using the Method of Lagrange Multipliers.


1
Expert's answer
2021-10-24T17:49:06-0400

We are asked for the extreme values of subject to the constraint "g(x, y)=x^2+y^2=1."

Using Lagrange multipliers, we solve the equations "\nabla f=\lambda \nabla g" and "g(x, y)=1", which can be written as


"f_x=\lambda g_x, f_y=\lambda g_y, g(x, y)=1"

or as


"4y=2\lambda x"

"4x=2\lambda y"

"x^2+y^2=1"

From two first equations

"\dfrac{y}{x}=\dfrac{x}{y}\ or\ x=y=0"

But "x^2+y^2=1." Then we have


"x^2=y^2=\dfrac{1}{2}"

Therefore has possible extreme values at the points "(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2}), (-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2}),"

"(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})," and "(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})." Evaluating at these four points, we find that


"f(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})=2"

"f(-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})=-2"

"f(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})=-2"

"f(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})=2"

Therefore the maximum value of on the circle is "f(-\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})=f(\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})=2" and the

minimum value is "f(-\dfrac{\sqrt{2}}{2}, \dfrac{\sqrt{2}}{2})=f(\dfrac{\sqrt{2}}{2}, -\dfrac{\sqrt{2}}{2})=-2."